Cary is 9 years older than dan. In 7 years the sum of their ages wil equall 93. Find both their ages now. Solve using equations.
Answer by algebrahouse.com(1659)

"Cary is 9 years older than dan. In 7 years the sum of their ages wil equall 93. Find both their ages now. Solve using equations."

x = Dan's age now
x + 9 = Cary's age now {Cary is 9 yrs older than Dan}

x + 7 = Dan's age in 7 years
x + 16 = Cary's age in 7 years

x + 7 + x + 16 = 93 {in seven years the sum of their ages will be 93}
2x + 23 = 93 {combined like terms}
2x = 70 {subtracted 23 from both sides}
x = 35 {divided both sides by 35}
x + 9 = 44 {substituted 35, in for x, into x + 9}

Dan is 35
Cary is 44
www.algebrahouse.com